OpenStudy (anonymous):

A theater charges $10 for children and $14 for adults to attend a drama. A total of 1600 tickets are sold for a total of &18400. How many children's ticket were sold?

OpenStudy (angela210793):

10x+14y=18400 whr x is the no.s of ticket which is sold to kids and y to adults x+y=1600(*-10) 10x+14y=18400 -10x-10y=-16000 (add both eq) 4y=2400--->y=600 so x+600=1600--->x=1000 children's ticket r sold

OpenStudy (amistre64):

ch@10 = (18400 - 1600(14))/(10-14) perhaps?

OpenStudy (amistre64):

that gives me 1000 ch ticks; dbl check tho

OpenStudy (angela210793):

my sol.isn't right? o.o

OpenStudy (angela210793):

ahhha :):)

OpenStudy (amistre64):

your way is the long stepped version; mine is simply a formula that condences it all into one shot

OpenStudy (angela210793):

tht formula doesn't seem familiar to me O.o idk...

OpenStudy (amistre64):

c + a = t cx + ay = T ; now eliminate the "a" -cy -ay = -ty cx +ay = T -------------- c(x-y) = T-ty c = (T-ty)/(x-y) c = (18400 - 1600(14)) / (10-14)

OpenStudy (angela210793):

hmmmm....interesting :) Thnx :)

OpenStudy (anonymous):

Thank you so much!!